OpenStudy (otonogold):

Multiple Choice Solve the exponential equation. (125^7x-2)=150 A) -0.1375 B) 2.1483 C) 0.4234 D) 0.4340

OpenStudy (anonymous):

do you know where to start?

OpenStudy (anonymous):

\[(125^{7x}-2)=150\] is this the question?

OpenStudy (otonogold):

take log from both sides & the 7x-2 is the exponent.

OpenStudy (otonogold):

I don't know what to do after I do the log on both sides..

OpenStudy (anonymous):

take log base 125 to each side

OpenStudy (anonymous):

start with that.. so you get 7x-2 on the left and log(150)/log(125) on the other side

OpenStudy (anonymous):

add 2 to each side.. so you get 7x=2+log(150)/log125)

OpenStudy (anonymous):

with me?

OpenStudy (otonogold):

yes

OpenStudy (anonymous):

solve for x.

OpenStudy (anonymous):

i should have typed ln instead of log .. sorry

OpenStudy (anonymous):

did you figure it out?

OpenStudy (agent0smith):

"take log base 125 to each side" if you do this, you'll have to use the change of base formula later... this is more work. Just take ln or log of both sides.

OpenStudy (anonymous):

sorry i meant to type that instead of log base ... haha sorry again

OpenStudy (otonogold):

You good lol I got = 0.4340

OpenStudy (anonymous):

x=2/7+(ln(150))/(7ln(125))

OpenStudy (anonymous):

b-i-n-g-0 !!

OpenStudy (anonymous):

good luck!

OpenStudy (otonogold):

(7x-2) * log(125) = log(150) divide both sides by log(125) 7x - 2 = log(150) / log(125) = 1.0378 Add 2 to both sides 7x = 3.0378 divide both sides by 7 x = 3.0378 / 7 = 0.4340 my work lol

OpenStudy (otonogold):

Thank you !(:

OpenStudy (agent0smith):

^very good work